2, 7, 26, 101, 400, ?
Answer by JulietG(1812)

* 4 -nth
.
(2 * 4) - 1 = 7
(7 * 4) - 2 = 26
(26 * 4) - 3 = 101
(101 * 4) - 4 = 400
(400 * 4) - 5 = ?